OpenStudy (juana02):

Two Coins. Flip two coins 50 times and record the number of times exactly one head was obtained. Determine the empirical probability of flipping exactly one head. How does this compare to the theoretical probability of flipping a coin and getting heads.

OpenStudy (anonymous):

Before I answer this one I wanted to make sure you know that the employee question answer was 190

OpenStudy (juana02):

Yes I'm having trouble with word problems I don't know how to substitute the words to solve them

OpenStudy (anonymous):

Do they give a number of times heads appeared?

OpenStudy (juana02):

It's says one head but no other given information

OpenStudy (kirbykirby):

I think they actually want you to flip two coins first to determine their empirical probability. THis would be calculated as: \[\frac{number ~of ~ times~ you~get~1~heads}{total~ number ~of ~flips,~i.e.~ 50 }\]

OpenStudy (juana02):

Ok I get dividing 1 into 50 0.002?

OpenStudy (kirbykirby):

Hmm well take a coin and flip it 50 times. You'll get tails or heads on some flips. So, Say you flip the coin 6 times, and you obtain: Tails, Heads, Heads, Heads, Tails, Heads. You get heads 4 times, out of 6 flips, so the empirical probability is 4/6 = 1/3 The theoretical probability on one flip is: \[\frac{number~ of~ ways ~to ~get ~heads}{number~of~poisible~outcomes}\] You can only obtain "heads" 1 way , and the number of possible outcomes is 2 (since you can get heads or tails)

OpenStudy (juana02):

I'm still confused @kirbykirby